OpenStudy (anonymous):

In a lottery there are 9 balls, numbered 1 to 9, in each of three bins. One ball will be selected from each bin. What are the odds of the numbers 2, 3, and 4 being pulled IN THAT ORDER?

OpenStudy (anonymous):

are the balls replaced?

OpenStudy (anonymous):

There is only 1 way to get 2,3,4 in that order. The total number of ways for the 3 balls is 9*8*7 = 504. So, success = 1, total = 504, fail = 503. odds = success / failure, so 1:503.

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

But they are in different bins .@PeachRings I am confused .. @akshay1234 that's not a choice.

OpenStudy (acxbox22):

you can pull a 2 out of 9 numbers so the chance is 1/9 you can then pull 3 out of 9 numbers the chance is again 1/9 finally you can pull 4 out of 9 numbers that is also 1/9 chance 1/9 x 1/9 x 1/9= 1/ 729

OpenStudy (anonymous):

728 is a choice but not 729 @acxbox22

OpenStudy (anonymous):

@david111 they are different bins . Would it still be1; 503 ?

OpenStudy (acxbox22):

are you sure?

OpenStudy (anonymous):

I am positive .

OpenStudy (anonymous):

acxbox should be right by my math

OpenStudy (anonymous):

whats the choices

OpenStudy (anonymous):

i believe its 1;503

OpenStudy (anonymous):

@david111 1 to 8 1 to 80 1 to 503 1 to 728

OpenStudy (anonymous):

Well my answer was on target so it's up to you

OpenStudy (anonymous):

a

OpenStudy (acxbox22):

i googled it and 1/503 is your answer but it looks like @david111 copied and pasted her answer http://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.250330.html here is proof no offence meant though

OpenStudy (anonymous):

so that means im right haha

OpenStudy (anonymous):

@david111 @acxbox22 I am about to post one other one . See if you agree with my answer I put on that one .